Please help me with these questions!. Will give thumbs up thank you!

0 0
Answer #1

IF YOU HAVE ANY DOUBTS COMMENT BELOW I WILL BE TTHERE TO HELP YOU

ANSWER:

EXPLANATION:

1)

Here the graph is

Number (Y) 350 300 250 200 150 100 50 Number (Y) 10 15 20 25 30

Here the cubic equation is

y ^ = -0.013x3 + 1.085x2 - 6.549x + 8.809

(b) Here the y^ (20000) = -0.013 * 153 + 1.085* 152  - 6.549 * 15 + 8.809 = 110.824

HEre the actual value = 109.5

so here the predicted value underestimated the actual value and it underestimate it by (110.824 - 109.5 = 1.824 )

(c) Here the actual population was = 300 million but there are more subscribers than the the population that means that on average there are more than 1 cell phone connection per capita in US. That could be possible due to multiple connections opted by the US general population.

HOPE IT HELPS YOU

RATE THUMBSUP IT HELPS ME ALOT

THANKS